OpenStudy (anonymous):

sec^2x/tanx = cotx+tanx I believe the first step is applying Pythagorean identity to get 1+tan^2x/tanx. Where do I go from here?

zepdrix (zepdrix):

\[\large\rm \frac{\sec^2x}{\tan x}\quad=\quad \frac{1+\tan^2x}{\tan x}\quad=\quad \frac{1}{\tan x}+\frac{\tan^2x}{\tan x}\]

zepdrix (zepdrix):

Applying this type of idea to break up the fraction:\[\large\rm \frac{a+b}{c}\quad=\quad \frac{a}{c}+\frac{b}{c}\]

OpenStudy (anonymous):

Oh ok, then the common factor is removed leaving 1/tanx + tan x

OpenStudy (anonymous):

Ohh ok I will write that down as well, the fraction breaking up. I never really understood that, thanks so much.

zepdrix (zepdrix):

np \c:/ gj